OpenStudy (anonymous):

The life expectancy of a typical lightbulb is normally distributed with a mean of 2,100 hours and a standard deviation of 40 hours. What is the probability that a lightbulb will last between 1,965 and 2,165 hours?

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@kropot72

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@camerondoherty

OpenStudy (kropot72):

First you need to find the z-scores for 1,965 and 2,165.

OpenStudy (anonymous):

How?

OpenStudy (kropot72):

\[\large z=\frac{X-\mu}{\sigma}\]

OpenStudy (anonymous):

What do the symbols mean again?

OpenStudy (anonymous):

I know z is z-score

OpenStudy (kropot72):

You should know this. How are you studying statistics?

OpenStudy (anonymous):

Hold on

OpenStudy (anonymous):

I am sorry, I just don't get this

OpenStudy (kropot72):

\[\large \mu=population\ mean\] \[\large \sigma=population\ standard\ deviation\]

OpenStudy (kropot72):

X = random variable

OpenStudy (anonymous):

So, The z-score=random variables-2100(population mean)/40(population standard deviation)

OpenStudy (anonymous):

What would the random variable be?

OpenStudy (kropot72):

There are two random variables. One is 1,965 and the other is 2,165. For the question, two fixed values of the random variable have been chosen.

OpenStudy (kropot72):

So we get the z-score for 1,965 as follows: \[\large z _{1}=\frac{1965-2100}{40}=you\ can\ calculate\]

OpenStudy (anonymous):

Hold on

OpenStudy (anonymous):

.004

OpenStudy (kropot72):

Not really. Lets try the calculation in steps. First the numerator: 1965 - 2100 = ?

OpenStudy (anonymous):

.0004

OpenStudy (kropot72):

Are you using a calculator?

OpenStudy (anonymous):

No, I just forgot to add 3 zero

OpenStudy (anonymous):

Not 2

OpenStudy (anonymous):

.0004

OpenStudy (kropot72):